# Tight Robustness Certificates and Wasserstein Distributional Attacks for Deep Neural Networks

### v1

@@ 3.1 Exact and tractable interpretation of WDRO for ReLU neural networks
where $J\in\partial\theta(x)$ is a general Jacobian of $\theta$ at $x$, then Jordan & Dimakis (2020, Theorem 1) has shown that $\|\theta(x^{\prime})-\theta(x)\|_{\tilde{r}}\leq L_{\theta}\times\|x^{\prime}-x\|_{r}$ for any $x^{\prime},x\in\mathcal{X}$. Moreover, if $\theta$ is in general position (Jordan & Dimakis, 2020, Definition 4), then the chain rule applies and any general Jacobian $J$ must has a form as $W_{H+1}D_{H}W_{H}\cdots D_{1}W_{1}$ for some $[0,1]$-diagonal matrix $D_{h}$, $h=1,\dots,H$. It is worth noting that the set of ReLU networks not in general position is negligible (Jordan & Dimakis, 2020, Theorem 3). Now in equation 5, the maximizer of a convex function (norm operator) must happen at vertices, thus we only need to consider $0/1$-diagonal matrix $D_{h}$.
@@
Given this definition of mask and note that $\mathcal{D}_{\mathcal{X}}$ is finite, one can rewrite equation 5 as $L_{\theta}=\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\|J_{\bm{D}}\|_{r\rightarrow\tilde{r}}$. We adopt this notion and show that it induces an upper bound for the Wasserstein distributional robust optimization (WDRO) problem 3 with cross-entropy loss. Moreover, this upper bound is tight for a class of monotonic ReLU networks.
Theorem 3.1 (WDRO for ReLU).
Given a ReLU network $\theta(x)=W_{H+1}(\operatorname{ReLU}(\cdots(W_{1}x+b_{1})\cdots)+b_{H})$ being in general position, $1/r+1/s=1$ and $\ell$ being the cross-entropy or DLR loss, define
$\bm{L}\triangleq 2^{1/s}\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\left\|J_{\bm{D}}\right\|_{{r}\to{s}},$ (6)
@@
Moreover, if the dual-norm maximizer $\mathcal{M}_{r}(J_{\bm{D}^{\star}}^{\top}(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}}))\in\mathrm{rec}({\mathcal{C}_{\bm{D}^{\star}}})$ where $\bm{D}^{\star}$ is a maximizer of 6 and $(k^{\prime\star},k^{\star})$ is a maximizer of 7, and $(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}})$ is the largest increment direction of $J_{\bm{D}^{\star}}$, then $\bm{l}=\bm{L}$.
@@
To prove inequality 8, we show that $\ell(\cdot,\theta)$ is $\bm{L}$-Lipschitz, and a direction $u$ found in equation 7 induces an admissible attack $\mathbb{P}_{\mathrm{adv}}$ satisfying that $\mathbb{E}_{\mathbb{P}_{\mathrm{adv}}}[\ell(Z;\theta)]\approx\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{l}\epsilon$ and $\mathcal{W}_{d,1}(\mathbb{P}_{\mathrm{adv}},\mathbb{P}_{N})\leq\epsilon$. To verify the sufficient condition of $\bm{l}=\bm{L}$, we show that the constructed $\mathbb{P}_{\mathrm{adv}}$ provides $\bm{l}=\bm{L}$. We provide detailed proof in Appendix A.1. ∎
In Figure 2, we illustrate an instance in which our lower and upper bounds match. While equation 7 provides a tight lower bound of the WDRO, it is impractical to scan through all $x\in\mathcal{X}$ and its mask $\mathcal{D}_{x}$. We then introduce a practical lower bound of which we consider the mask associated with the sample points only.
Figure 2: WDRO bounds and PGD attack loss for a fixed $n=K=2$ ReLU classifier with one hidden layer of dimension 8. Lower-bound curves are the cumulative $\bm{l}$ as more reachable activation masks are considered.
@@
$\bm{l}_{N}\triangleq\max_{\begin{subarray}{c}(X^{(i)},Y^{(i)})\in\mathcal{Z}_{N},\\ \bm{D}\in\mathcal{D}_{x}\end{subarray}}\max_{k}\sup_{\|u\|_{r}=1}\left\{(\bm{e}_{k}-Y^{(i)})^{\top}J_{\bm{D}}u\mid u\in\mathrm{int}(\mathrm{rec}({\mathcal{C}_{\bm{D}}}))\right\}.$ (9)
@@
In the spirit of formulation 9, we propose a practical adversarial attack in Section 4 which aims to find adversarial direction $u^{(i)}$ for each sample $i$ so that it maximizes the change of the logit function.
@@ 3.2 Exact and tractable interpretation of WDRO for smooth activation neural networks
For networks with smooth activations, e.g, GELU (Hendrycks & Gimpel, 2016), SiLU/Swish (Ramachandran et al., 2017; Elfwing et al., 2018), WDRO duality connects worst-case (adversarial) risk to first-order geometry via the Jacobian of the logit map, yielding global Lipschitz-type upper penalties of the form $\sup_{x\in\mathcal{X}}\|J(x)\|_{r\to s}$. Compared to piecewise-linear ReLU certificates, smooth nets trade exact cell-wise constancy for differentiability along rays and curves, suggesting bounds driven by asymptotic Jacobian behavior rather than activation masks.
Let $t$ be a positive scalar, $\theta:\mathcal{X}\subseteq\mathbb{R}^{n}\to\mathbb{R}^{K}$ be a classifier with smooth activations and cross-entropy loss; let $J(x)\in\mathbb{R}^{K\times n}$ be its Jacobian. We then have the following result.
@@
Let $\theta:\mathbb{R}^{n}\rightarrow\mathbb{R}^{k}$ be a differentiable network, $1/r+1/s=1$ and $\ell$ being the cross-entropy or DLR loss, define
$\bm{L}\triangleq{2^{1/{s}}}\sup_{x\in{\mathcal{X}}}\left\|\nabla_{x}\theta(x)\right\|_{{r}\to{s}},$
@@
$\bm{l}\triangleq\sup_{x\in{\mathcal{X}}}\max_{k^{\prime}\neq k}\sup_{\|u\|_{r}=1}\left\{(\bm{e}_{k^{\prime}}-\bm{e}_{k})^{\top}\nabla_{x}\theta(x)u\right\}.$
@@
In this setting, first-order WDRO penalties are controlled by how $J(x)$ amplifies unit directions and how that amplification projects onto the most competitive non-true logit. The upper slope $\bm{L}$ is the global worst-case amplification, while the lower, margin–directional slope $\bm{l}$ follows rays $x^{(i)}+tu$ and harvests only the component along $(\bm{e}_{k^{\prime}}-\bm{e}_{k_{i}})$. When a ray both attains the global operator norm and aligns with a margin difference, the bound is tight to first order ($\bm{l}=\bm{L}$). This motivates the adversarial procedure used our adversarial attack algorithm WDA (Algogirthm 1), where we search for a direction $u$ and a rival class $k^{\prime}$ that maximize the first-order increase $(\bm{e}_{k^{\prime}}-\bm{e}_{k_{i}})^{\top}J(x)u$.
4 Wasserstein Distributional Attack
@@
Existing point-wise attacks such as FGSM (Goodfellow et al., 2014), DeepFool (Moosavi-Dezfooli et al., 2016), AA (Croce & Hein, 2020), AAA (Liu et al., 2022), keep the adversarial distribution supported on exactly $N$ points, where each point $X^{(i)}_{\operatorname{adv}}$ is perturbed to be precisely on the boundary of the $\epsilon$-ball centered at $X^{(i)}$. To address this issue, we propose a novel method called Wasserstein Distributional Attack (WDA). At a high level, WDA constructs an adversarial distribution, $\mathbb{P}_{\mathrm{adv}}$, supported on a set of $2N$ points. This set comprises of $N$ original empirical samples $X^{(i)}$ and $N$ corresponding adversarial points $X^{(i)}_{\mathrm{adv}}$, each perturbed to an $r$-norm distance $\kappa\epsilon$ from $X^{(i)}$ using the first-order, margin-aligned directions predicted by Theorems 3.13.3, for some $\kappa\geq 1$. In other words,
$\mathbb{P}_{\mathrm{adv}}=\frac{1}{N}\sum_{i=1}^{N}\left(1-\frac{1}{\kappa}\right)\bm{\delta}_{(X^{(i)},Y^{(i)})}+\frac{1}{\kappa}\bm{\delta}_{(X_{\mathrm{adv}}^{(i)},Y^{(i)})}.$ (10)
In the special case where $\kappa=1$, our proposed attack reduces to existing point-wise methods. When $\kappa=2$, WDA simplifies to a uniform distribution over all $2N$ points, with each point receiving a weight of $\frac{1}{2N}$. This $2N$-support mixture belongs to the $\Omega_{1}$ ambiguity set and serves as a constructive, distributional adversary; it is not necessarily the inner maximizer of WDRO. We now make the first-order ascent directions explicit; this is the step used by WDA to realize the margin-aligned rays from Theorems 3.13.3.
Define (sub)gradient $g_{j}(x)\in\partial_{x}\big(\theta_{j}-\theta_{k}\big)(x)$. Then $\mathcal{M}_{r}(g_{j})$ give the per–iteration, first–order version of the ray ascent used in Theorems 3.13.3: within a ReLU cell (affine logits) or for smooth activations (continuous $J$), moving along $u_{j}=\mathcal{M}_{r}(g_{j})$ increases the gap at rate $\|g_{j}(x)\|_{s}$. During an initial probing phase, we evaluate all rivals $j\neq k$ using these first-order steps. At the end of that phase, we fix a single rival $j^{*}$ based on the logits magnitude and continue the remaining iterations. If we allow $j^{*}$ to change at every step, the update can oscillate across classes and chase locally steep but globally suboptimal directions for misclassifications. Finally, we project each step to the ball of radius $\kappa\varepsilon$ around the anchor $X^{(i)}$ to the WDRO budget. The procedure for implementing the Wasserstein Distributional Attack is presented in Algorithm 1. A visualization of our algorithm is shown in Figure 3.
@@
1:Inputs: neural network $\theta:\mathbb{R}^{n}\to\mathbb{R}^{K}$, empirical distribution $\mathbb{P}_{N}=\sum_{i=1}^{N}\bm{\delta}_{(X^{(i)},Y^{(i)})}$, budget $\epsilon>0$, cost-norm $r\in\{1,2,\infty\}$, WDA paramter $\kappa\geq 1$, step size $\alpha>0$, and $0<$ prob $\leq$ maxiter
@@
2:Outputs: Wasserstein distributional attack $\mathbb{P}_{\mathrm{adv}}$ such that $\mathcal{W}_{d,1}(\mathbb{P}_{\mathrm{adv}},\mathbb{P}_{N})\leq\epsilon$ where $d((x^{\prime},y^{\prime}),(x,y))=\|x^{\prime}-x\|_{r}+\infty\cdot\bm{1}_{\{y^{\prime}\neq y\}}$
3:Initialize: dual-norm maximizer $\mathcal{M}$ (1), projection $\Pi$ (2)
4:for $i=1$ to NN do
5:  $x_{0}\leftarrow X^{(i)}$, $\bm{e}_{k}\leftarrow Y^{(i)}$ for some $k=1,\dots,K$
6:  for iter =0=0 to maxiter do
7:   if iter << prob then $\mathcal{J}=\{1,\dots,K\}\setminus\{k\}$ else $\mathcal{J}=\{j^{*}\}$
@@
8:   $g_{j}\leftarrow\nabla_{x}\theta\left(x_{\mathrm{iter}}\right)^{\top}(\bm{e}_{j}-\bm{e}_{k})$ for $j\in\mathcal{J}$
9:   $u_{j}\leftarrow\mathcal{M}_{r}(g_{j})$ for $j\in\mathcal{J}$
@@
10:   $\varphi_{j}\leftarrow\Pi_{r,X^{(i)},\kappa\epsilon}\left(x_{\mathrm{iter}}+\alpha u_{j}\right)$ for $j\in\mathcal{J}$
11:   $j^{*}=\arg\max_{j\in\mathcal{J}}\theta_{j}\left(\varphi_{j}\right)$
12:   $x_{\mathrm{iter+1}}\leftarrow\varphi_{j^{*}}$
13:  end for
14:  $X^{(i)}_{\mathrm{adv}}\leftarrow x_{\mathrm{maxiter}}$
15:end for

### v2

@@ 3.1 Exact and tractable interpretation of WDRO for ReLU neural networks
where $J\in\partial\theta(x)$ is a general Jacobian of $\theta$ at $x$, then Jordan and Dimakis (2020, Theorem 1) has shown the inequality that $\|\theta(x^{\prime})-\theta(x)\|_{\tilde{r}}\leq L_{\theta}\times\|x^{\prime}-x\|_{r}$ for any $x^{\prime},x\in\mathcal{X}$. Moreover, if $\theta$ is in general position (Jordan and Dimakis, 2020, Definition 4), then the chain rule applies and any general Jacobian $J$ must have a form as $W_{H+1}D_{H}W_{H}\cdots D_{1}W_{1}$ for some $[0,1]$-diagonal matrix $D_{h}$, $h=1,\dots,H$. It is worth noting that the set of ReLU networks not in general position is negligible (Jordan and Dimakis, 2020, Theorem 3). Now in equation 5, the maximization of a convex function (norm operator) is achieved at the vertices, thus we only need to consider $0/1$-diagonal matrix $D_{h}$.
@@
Given this definition, and noting that $\mathcal{D}_{\mathcal{X}}$ is finite, one can rewrite equation 5 as $L_{\theta}=\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\|J_{\bm{D}}\|_{r\rightarrow\tilde{r}}$. We adopt this notion and show that it induces an upper bound for the Wasserstein distributional robust optimization (WDRO) problem equation 3 with cross-entropy loss. Moreover, this upper bound is tight for a class of monotonic ReLU networks.
Theorem 3.1 (WDRO for ReLU Networks).
Given a ReLU network
$\theta(x)=W_{H+1}(\operatorname{ReLU}(\cdots(W_{1}x+b_{1})\cdots)+b_{H})$
being in general position, $1/r+1/s=1$ and $\ell$ being the cross-entropy or DLR loss, define
$\bm{L}\triangleq 2^{1/r}\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\left\|J_{\bm{D}}\right\|_{{r}\to{s}},$ (6)
@@
Moreover, let $\bm{D}^{\star}$ be a maximizer of equation 6, $(k^{\prime\star},k^{\star})$ be a maximizer of equation 7 and $(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}})$ be the largest increment direction of $J_{\bm{D}^{\star}}$. If the dual-norm maximizer $\mathcal{M}_{r}(J_{\bm{D}^{\star}}^{\top}(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}}))\in\mathrm{rec}({\mathcal{C}_{\bm{D}^{\star}}})$ then $\bm{l}=\bm{L}$.
@@
To prove inequality equation 8, we show that the loss function $\ell(\cdot,\theta)$ is $\bm{L}$-Lipschitz, and a direction $u$ found in equation 7 induces an admissible attack distribution $\mathbb{P}_{\mathrm{adv}}$ satisfying $\mathbb{E}_{\mathbb{P}_{\mathrm{adv}}}[\ell(Z;\theta)]\approx\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{l}\epsilon$, $\mathcal{W}_{d,1}(\mathbb{P}_{\mathrm{adv}},\mathbb{P}_{N})\leq\epsilon$. To verify the sufficient condition of $\bm{l}=\bm{L}$, we show that the constructed $\mathbb{P}_{\mathrm{adv}}$ provides $\bm{l}=\bm{L}$. We provide detailed proof in Appendix B.1. ∎
In Figure 1, we illustrate an instance in which our lower and upper bounds match. While equation 7 provides a tight lower bound of the WDRO, it is impractical to scan through all $x\in\mathcal{X}$ and its mask $\mathcal{D}_{x}$. We then introduce a practical lower bound, of which we consider the mask associated with the sample points only.
@@
$\bm{l}_{N}\triangleq\max_{\begin{subarray}{c}(X^{(i)},Y^{(i)})\in\mathcal{Z}_{N}\\ \bm{D}\in\mathcal{D}_{x}\end{subarray}}\max_{k}\sup_{\|u\|_{r}=1}\left\{(\bm{e}_{k}-Y^{(i)})^{\top}J_{\bm{D}}u\mid u\in\mathrm{int}(\mathrm{rec}({\mathcal{C}_{\bm{D}}}))\right\}.$ (9)
@@
Based on the proof of our lower bound equation 17, we construct a worst-case distribution by moving mass from a sample along a direction $u$ that maximizes the margin term in equation 7. In Section 4, based on formulation equation 9, we create this construction empirically via the attack distribution equation 11 by choosing adversarial direction $u^{(i)}$ for each sample $i$ so that it maximizes the first-order increase of the corresponding logit margin.
@@ 3.2 Exact and tractable interpretation of WDRO for smooth activation neural networks
While ReLU networks require analyses of activation masks, modern architectures often employ smooth activation functions (e.g., GELU (Hendrycks and Gimpel, 2016), SiLU (Ramachandran et al., 2017)). In these settings, the WDRO duality connects the worst-case adversarial risk directly to the first-order geometry via the Jacobian of the logit map.
@@
Given a differentiable network $\theta:\mathbb{R}^{n}\rightarrow\mathbb{R}^{K}$ with $J_{\theta}(x)\in\mathbb{R}^{K\times n}$ being its Jacobian, $1/r+1/s=1$ and $\ell$ being the cross-entropy or DLR loss, define
$\bm{L}\triangleq 2^{1/r}\sup_{x\in{\mathcal{X}}}\left\|J_{\theta}(x)\right\|_{{r}\to{s}},$
@@
$\bm{l}\triangleq\sup_{x\in{\mathcal{X}}}\max_{k^{\prime}\neq k}\sup_{\|u\|_{r}=1}(\bm{e}_{k^{\prime}}-\bm{e}_{k})^{\top}J_{\theta}(x)u.$
@@
In addition, this result extends certification guarantees to modern differentiable Transformer architectures such as GPT (Radford et al., 2019), ViT (Dosovitskiy et al., 2021), and Swin Transformer (Liu et al., 2021), which are often Pre-LN Transformers Xiong et al. (2020).
Corollary 3.4 (WDRO for Pre-LN Transformers).
Given assumptions and notations used in Theorem 3.3, if $\theta$ is a depth-$H$ (Pre-LN) Transformer, then by Castin et al. (2024, Theorem 3.3 and Lemma 3.8), the upper bound satisfies
$\bm{L}=2^{1/r}\sup_{x\in\mathcal{X}}\|J_{\theta}(x)\|_{r\to s}\leq 2^{1/r}(nK)^{(\frac{1}{2}-\frac{1}{r})_{+}}\times\Lambda_{\theta},$ (10)
where $\Lambda_{\theta}\triangleq L_{\rm head}L_{\rm emb}\prod_{h=1}^{H}\left(1+L_{\rm LN}^{(h)}L_{\rm MHA}^{(h)}\right)\left(1+L_{\rm LN^{\prime}}^{(h)}L_{\rm FFN}^{(h)}\right)$, and all module constants are Euclidean Lipschitz constants on $\mathcal{X}$. Moreover, their explicit bound of the form $L_{\rm MHA}^{(h)}=\mathcal{O}(\sqrt{n})$ implies $\bm{L}=\mathcal{O}\left(2^{1/r}(nK)^{(\frac{1}{2}-\frac{1}{r})_{+}}\times n^{H/2}\right)$ up to architecture and parameter-dependent constants.
4 Wasserstein Distributional Attacks
@@
Existing point-wise attacks such as FGSM (Goodfellow et al., 2014), DeepFool (Moosavi-Dezfooli et al., 2016), AA (Croce and Hein, 2020), AAA (Liu et al., 2022), keep the adversarial distribution supported on exactly $N$ points, where each point $X^{(i)}_{\operatorname{adv}}$ is perturbed to be precisely on the boundary of the $\epsilon$-ball centred at $X^{(i)}$. To address this issue, we propose a novel method called Wasserstein Distributional Attacks which constructs an adversarial distribution $\mathbb{P}_{\mathrm{adv}}$, supported on a set of $2N$ label-preserving points
$\mathbb{P}_{\mathrm{adv}}=\frac{1}{N}\sum_{i=1}^{N}\left(1-\frac{1}{\kappa_{i}}\right)\bm{\delta}_{(X^{(i)},Y^{(i)})}+\frac{1}{\kappa_{i}}\bm{\delta}_{(X^{(i)}_{\mathrm{adv}},Y^{(i)})},$ (11)
where $\kappa_{i}\geq 1$. (See Figure 2.) This $2N$-support consists of $N$ original empirical samples $X^{(i)}$ and $N$ corresponding adversarial points $X^{(i)}_{\mathrm{adv}}$, each perturbed using the first-order, margin-aligned directions predicted by Theorems 3.1 and 3.3.
WDA.
WDA corresponds to $\kappa_{i}\equiv\kappa$. In the special case where $\kappa=1$, the attack reduces to point-wise methods, while $\kappa=2$ yields a uniform distribution over all $2N$ points, with each point receiving a weight of $\frac{1}{2N}$. The mixture equation 11 lies in $\Omega_{p}$ whenever $\|X^{(i)}_{\mathrm{adv}}-X^{(i)}\|_{r}\leq\kappa^{1/p}\epsilon$ for all $i$ (Corollary C.2) and serves as a constructive and feasible distributional adversary.
WDA++.
WDA++ keeps the same two-point anchored form equation 11 but chooses the mixing weight adaptively. For each sample $i$ we calculate $d_{i}$, an estimate of the closest distance to flip an image’s label and set $1/\kappa_{i}=\min\{1,t_{i}/d_{i}^{p}\}$ for a transport allocation $t_{i}\geq 0$ (with the conventions $1/\kappa_{i}=1$ if $d_{i}=0$ and $1/\kappa_{i}=0$ if $d_{i}=\infty$), yielding $X^{(i)}_{\rm adv}$ chosen near the closest flip. By Corollary C.2, $\mathcal{W}_{d,p}(\mathbb{P}_{\rm adv},\mathbb{P}_{N})^{p}\leq\sum_{i}\mu_{i}t_{i}$, hence $\mathbb{P}_{\rm adv}\in\Omega_{p}$ whenever $\sum_{i}\mu_{i}t_{i}\leq\epsilon^{p}$. Full implementation details are deferred to Appendix E.
Figure 3: Wasserstein Distributional Attack (WDA, Algorithm 1) for $r=2$. At each iteration $x_{t}$, WDA forms K1K\!-\!1 candidates $\varphi_{j}$ and updates using the one with the largest logit $\theta_{j}(\varphi_{j})$. For reference, PGD follows the dual-norm gradient direction; DeepFool linearizes the decision boundary.
@@
Inputs: neural network $\theta:\mathbb{R}^{n}\to\mathbb{R}^{K}$, empirical distribution $\mathbb{P}_{N}=\frac{1}{N}\sum_{i=1}^{N}\bm{\delta}_{(X^{(i)},Y^{(i)})}$, budget $\epsilon>0$, cost-norm $r\in\{1,2,\infty\}$, Wasserstein order $p\in\{1,2\}$, WDA parameter $\kappa\geq 1$, step size $\alpha>0$, and $0<$ probe $\leq$ maxiter
@@
Outputs: Wasserstein distributional attack $\mathbb{P}_{\mathrm{adv}}$ such that $\mathcal{W}_{d,p}(\mathbb{P}_{\mathrm{adv}},\mathbb{P}_{N})\leq\epsilon$ where $d((x^{\prime},y^{\prime}),(x,y))=\|x^{\prime}-x\|_{r}+\infty\cdot\bm{1}_{\{y^{\prime}\neq y\}}$
Initialize: dual-norm maximizer $\mathcal{M}$ equation 1, projection $\Pi$ equation 2
for $i=1$ to NN do
  $x_{0}\leftarrow X^{(i)}$, $\bm{e}_{k}\leftarrow Y^{(i)}$ for some $k=1,\dots,K$
  for iter =0=0 to maxiter do
   if iter << probe then $\mathcal{J}=\{1,\dots,K\}\setminus\{k\}$ else $\mathcal{J}=\{j^{*}\}$
@@
   $g_{j}\leftarrow\nabla_{x}\theta\left(x_{\mathrm{iter}}\right)^{\top}(\bm{e}_{j}-\bm{e}_{k})$ for $j\in\mathcal{J}$
   $u_{j}\leftarrow\mathcal{M}_{r}(g_{j})$ for $j\in\mathcal{J}$
@@
   $\varphi_{j}\leftarrow\Pi_{r,X^{(i)},\kappa^{1/p}\epsilon}\left(x_{\mathrm{iter}}+\alpha u_{j}\right)$ for $j\in\mathcal{J}$
   $j^{*}\leftarrow\arg\max_{j\in\mathcal{J}_{i}}\left(\theta_{j}(\varphi_{j})-\theta_{k}(\varphi_{j})\right)$
   $x_{\mathrm{iter+1}}\leftarrow\varphi_{j^{*}}$
  end for
  $X^{(i)}_{\mathrm{adv}}\leftarrow x_{\mathrm{\texttt{maxiter}}}$
end for
$\mathbb{P}_{\mathrm{adv}}\leftarrow\frac{1}{N}\sum_{i=1}^{N}\left(1-\frac{1}{\kappa}\right)\bm{\delta}_{(X^{(i)},Y^{(i)})}+\frac{1}{\kappa}\bm{\delta}_{(X_{\mathrm{adv}}^{(i)},Y^{(i)})}$
return $\mathbb{P}_{\mathrm{adv}}$
Adversarial direction.
Define the (sub)gradient $g_{j}(x)\in\partial_{x}\big(\theta_{j}-\theta_{k}\big)(x)$. Then $\mathcal{M}_{r}(g_{j})$ provides the per-iteration, first-order instance of the ray ascent from Theorems 3.1 and 3.3: within a ReLU cell or for smooth activations, moving along $u_{j}=\mathcal{M}_{r}(g_{j})$ increases the gap at rate $\|g_{j}(x)\|_{s}$. During an initial probing phase, we evaluate all rivals $j\neq k$ using these first-order steps. At the end of this phase, we fix a single rival $j^{*}$ based on the logit magnitude and continue the remaining iterations. Fixating $j^{*}$ prevents the update from oscillating across classes or chasing locally steep but globally suboptimal directions. Finally, we project each step to the ball of radius $\kappa^{1/p}\varepsilon$ around the anchor $X^{(i)}$. A visualization of our algorithm is shown in Figure 3. The procedures for implementing WDA and WDA++ are presented in Algorithms 1 and 2.
5 Related Work
Robustness certificates.
Early scalable global certificates control the Lipschitz constant by multiplying per-layer operator norms, which is fast to compute yet data-agnostic and typically loose on deep nets (Virmaux and Scaman, 2018). For ReLU networks, local (activation-aware) methods exploit piecewise linearity to produce much tighter, input-conditioned certificates on individual activation regions (Katz et al., 2017; Ehlers, 2017; Weng et al., 2018; Singh et al., 2018; Shi et al., 2022). Most relevant to exact local Lipschitzness, Jordan and Dimakis (2020) showed that for a broad class of ReLU networks in general position, the local Lipschitz constant can be computed exactly by optimizing over activation patterns.
Adversarial attacks.
Adversarial Attack methods seek for perturbation $x^{\prime}$ formed by adding a small, human-imperceptible perturbation to a clean input $x$ that causes misclassification (Szegedy et al., 2014). The threat model specifies the attacker’s knowledge (white-box vs. black-box), the admissible perturbation set (e.g., $r_{2}$ balls with budget $\epsilon$), and the objective (e.g., worst-case loss within the ball). Canonical white-box methods include FGSM (Goodfellow et al., 2014), multistep PGD (Madry et al., 2018), CW (Carlini and Wagner, 2017), and gradient-based margin attacks such as DeepFool (Moosavi-Dezfooli et al., 2016). In the context of distributional threats, Bai et al. (2023) introduced W-PGD which transports the empirical measure while maintaining its support size. This effectively restricts the attack to a point-wise one. Our work adopts a different perspective by expanding the support to $2N$ points, thereby exploring a larger region of the ambiguity set. Decision-based and score-free attacks (black-box) include Boundary Attack (Brendel et al., 2021) and Square Attack (Andriushchenko et al., 2020).
Robust evaluation is subtle: gradient masking can inflate apparent robustness if attacks are not adapted (Athalye et al., 2018). To standardize evaluation, AutoAttack (AA) (Croce and Hein, 2020) composes strong, parameter-free attacks (APGD-CE, APGD-DLR, FAB, Square) and is widely adopted for reporting robust accuracy. RobustBench (Croce et al., 2021) curates model zoos and standardized test protocols across datasets and $r_{p}$ threat models, enabling comparable and reproducible robustness claims. Liu et al. (2022) proposed Adaptive Auto Attack (A3), which incorporates Adaptive Direction Initialization (ADI) and Online Statistics-based Discarding (ODS) (Tashiro et al., 2020) with a PGD framework to enhance attack efficiency. In our experiments, we report robustness under AA and A3 following RobustBench conventions and use them as baselines for comparison.
Several works have focused on adversarial attacks tailored to ReLU networks. Croce and Hein (2018) introduced rLR-QP, a gradient-free method that navigates the piecewise-linear regions of ReLU models by solving convex subproblems and enhancing exploration with randomization and local search. More recently, Zhang et al. (2022) developed BaB-Attack, a branch-and-bound framework that operates in activation space, leveraging bound propagation, beam search, and large neighborhood search to uncover stronger adversarial examples than conventional gradient-based approaches, particularly on hard-to-attack inputs. As pointed out in Zhang et al. (2022); Croce et al. (2020), rLR-QP and BaB-Attack are not as efficient as gradient based attack, therefore, for single method ($\Omega_{\infty}$) baseline we only use APGD.
6 Comparison with Existing Baselines
Experimental setup.
We evaluate the effectiveness of WDA and WDA++ in amplifying misclassification rates against state-of-the-art defences on CIFAR-10, CIFAR-100, and ImageNet. We report robustness under $r_{\infty}$ and $r_{2}$ perturbations with budgets $\epsilon\in\{4/255,8/255,0.5\}$. Defence models and pre-trained weights are sourced from RobustBench (Croce et al., 2021). Our benchmarks include single-method baselines (APGD-CE, APGD-DLR, W-PGD) and ensemble methods (AA, A3).
For baselines, we copy the official default configurations.
For WDA, we use probe steps $\texttt{probe}=10$, attack iterations $\texttt{maxiter}=20$ and step size $\alpha=0.2$ for $r_{2},\epsilon=0.5$, $\alpha\in\{0.01,0.02\}$ for $r_{\infty},\epsilon\in\{4/255,8/255\}$ respectively. We further propose A3-WDA, an extension that substitutes the PGD attack of A3 with our WDA attack. In addition to standard point-wise evaluation with $\kappa=1$, we validate the distributional threat model using $\kappa=2$. We report the classification accuracy on the constructed distribution $\mathbb{P}_{\mathrm{adv}}$ as $(1-1/\kappa)\times\mathrm{acc_{clean}}+(1/\kappa)\times\mathrm{acc_{adv}}$. Comprehensive ablation studies are detailed in Appendix D.
For WDA++, we evaluate under the constrained $\Omega_{2}$ set (comparable to W-PGD) and the full Wasserstein ambiguity ball $\Omega_{1}$. We adopt the parameter settings of WDA and detail the implementation in Appendix E.
All experiments are conducted on 2x NVIDIA GeForce RTX 4090 and 1x NVIDIA H200 GPUs.
Results on CIFAR-10.
Table 1 presents performance on CIFAR-10. Under both $r_{\infty}$ and $r_{2}$ threat models, WDA++ ($\Omega_{1}$) significantly reduces robust accuracy compared to all other methods. This confirms that standard robustness benchmarks significantly overestimate true robustness against distributional shifts. Moreover, expanding the ambiguity set to the full Wasserstein ball ($\Omega_{1}$) exposes vulnerabilities that both distributional adversaries under $\Omega_{2}$ and point-wise attacks under $\Omega_{\infty}$ fail to detect. Within the $\Omega_{2}$ constraint, WDA++ produces lower robust accuracy compared to W-PGD, confirming that our attack finds better directions in the loss landscape and utilises the Wasserstein budget $\epsilon$ more efficiently. Within the point-wise setting ($\kappa=1$), WDA remains competitive, outperforming single-method attacks (APGD-CE and APGD-DLR) and achieves results comparable to ensemble-based methods, indicating its ability to match the strength of more computationally demanding attack aggregations. When integrated into the ensemble framework (A3-WDA), our method consistently achieves the lowest robust accuracy, surpassing the standard A3 and AA benchmarks.
Table 1: Comparison of robust accuracy of WDA and baseline methods against various defences on CIFAR-10, CIFAR-100 and ImageNet. The best (lowest) robust accuracy of single methods and ensemble methods are highlighted in underline and bold, respectively.